# Difference-in-Differences Designs: A Practitioner’s Guide

### v1

@@ 5.2.5 Estimators for staggered designs with covariates
As the discussions in Section 5.2.2 made it clear, we can view the staggered DiD setups as a collection of simpler $2\times 2$ DiD building blocks. A benefit of this interpretation is that, if a parallel trends is only plausible after conditioning on covariates that are determinants of the changes in untreated potential outcomes, we can straightforwardly leverage all the results discussed in Section 4 and identify, estimate and make inference about the $ATT(g,t)$’s using regression adjustment, inverse probability weighting, or doubly robust methods.
Of course, to proceed in this manner, we would need to adopt conditioned on covariates versions of Assumption PT-GT-Nev, PT-GT-NYT or PT-GT-all, as well as impose an overlap condition. Given that all these are fairly similar to each other, here we only state an extension of Assumption PT-GT-NYT and a strong overlap condition that can be used for all three cases.
@@
For every eventually treated group $g$, not-yet-treated group $g^{\prime}$ and time periods $t$ such that $t\geq g$ and $g^{\prime}>t$, and every covariate value $X_{i}$
@@
For every group $g\in\mathcal{G}$, the conditional (weighted) probability of belonging to a treatment group $g$, given observed covariates $X_{i}$ that are determinants of untreated potential outcome growth, is uniformly bounded away from zero and one. That is, for some $\epsilon>0$ and for every group $g\in\mathcal{G}$ and, $\epsilon<P_{\omega}[G_{i}=g|X_{i}]<1-\epsilon$.
Based on these identifying assumptions and building on the results in Sections 4 and 5.2.2, we can follow the arguments in Callaway2021 and establish that the post-treatment $ATT(g,t)$’s are identified by the regression adjusted, inverse probability weighted, and doubly robust estimands given by [^32] Wooldridge2021 propose alternative estimators for the $ATT(g,t)$ that incorporate covariates and their interactions with group and time dummies into (5.12). Although Borusyak2023 and deChaisemartin2020 also allow for covariates in their estimation procedure, they affect the outcome levels and not their changes over time. Thus, to some extent, these procedures do not build on conditional parallel trends like Assumption CPT-GT-NYT.
@@
Estimating the $ATT(g,t)$s follows exactly as in Section 4, and event study aggregations follow from the arguments in Section 5.2.4. Figure 9 reports event study summary estimates incorporating covariates into the Medicaid analysis. Similarly to Figure 8, the results also suggest no effect of Medicaid on adult mortality among counties that expanded Medicaid by 2019. Given the point estimates and uniform confidence interval, we can be reasonably confident that the treatment effects are not greater than 6 or less than 5 deaths per 100,000 adults for the six-year period following Medicaid expansion.
@@
Notes: This figure shows the event study estimates with staggered treatment timing using the doubly-robust estimation method from Callaway2021. The outcome variable is the crude mortality rate for adults ages 20-64, and the covariates include the percentage of the county population that is female, the percentage of the county population that is white, the percentage of the county population that is Hispanic, the unemployment rate, the poverty rate, and county-level median income. The point estimate is reported by the circles, and both 95% point-wise (black) and simultaneous (red) confidence intervals are reported with the vertical lines. We also report the simple average of all non-negative event times as a summary of the overall ATT (together with their standard errors and 95% confidence interval).
@@ 5.3 Limitations of TWFE regressions
Our framework emphasizes building an estimator from $2\times 2$ components each of which targets a well-defined $ATT$ parameter, under a specific parallel trends assumption. The most common estimator for staggered designs, a TWFE regression, comes instead from extending convenient estimation tools that work well in the $2\times 2$ case. A TWFE specification that estimates a summary treatment effect parameter is:
@@
To isolate this issue, consider a setting with two time periods and three groups: a group that enters treatment in the first period ($G_{i}=1$), a group that becomes treated in the second time period ($G_{i}=2$), and a never treated group ($G_{i}=0$). This is a staggered design because group 1 and group 2 are treated at different times, but because there are only two time periods, we can re-write the TWFE specification as
$\Delta Y_{i,2}=\Delta\theta_{t}+\beta^{twfe}\Delta D_{i,2}+\Delta e_{i,2}$
Because $\Delta D_{i,2}$ only takes two values—1 for units whose treatment status increases ($G_{i}=t$), and 0 for units whose treatment status does not change ($G_{i}=t-1$ and $G_{i}=0$)—the TWFE estimand is the following simple comparison of means:
$\beta^{twfe}$ $=$ $\mathop{}\!\mathbb{E}[\Delta Y_{i,2}|\Delta D_{i,2}=1]-\mathop{}\!\mathbb{E}[\Delta Y_{i,2}|\Delta D_{i,2}=0]$
$=$ $\Big{(}\mathop{}\!\mathbb{E}[\Delta Y_{i,2}|G_{i}=2]-\mathop{}\!\mathbb{E}[\Delta Y_{i,2}|G_{i}=0]\Big{)}(1-w_{1})+\Big{(}\mathop{}\!\mathbb{E}[\Delta Y_{i,2}|G_{i}=2]-\mathop{}\!\mathbb{E}[\Delta Y_{i,2}|G_{i}=2]\Big{)}w_{1}$
where $w_{1}=\frac{p_{1}}{p_{1}+p_{0}}$ and $p_{g}=P(G=g)$ is group $g$’s share of units. The TWFE coefficient, in this case, is a weighted average of two DiD terms that use the already-treated units or the never-treated units as comparisons. We have already discussed how under PT between group $G_{i}=2$ and never-treated units, the first term in $\widehat{\beta}^{twfe}$ equals $ATT(2,2)$. But what about the second term with the already-treated comparison group? It turns out that by subtracting a trend in treated outcomes, this type of estimator generally equals a combination of treatment effects for both groups:
$\mathop{}\!\mathbb{E}[\Delta Y_{i,2}|G_{i}=2]-\mathop{}\!\mathbb{E}[\Delta Y_{i,2}|G_{i}=2]=ATT(2,2)-\Big{(}ATT(1,2)-ATT(1,1)\Big{)}\par$ (5.17)
@@
$\beta^{twfe}=ATT(2,2)+ATT(1,1)w_{1}-ATT(1,2)w_{1}$ (5.18)
The $ATT(1,2)$ building block receives negative weight in the overall TWFE estimand unless there are no treatment effect dynamics ($ATT(2,1)=ATT(2,2)$). In this example, the problem is easy to fix: drop the always-treated units and target $ATT(2,2)$. With multiple periods, however, the problem is more complex.
There are two primary results regarding TWFE estimators for general staggered timing designs. Goodman2021 decomposes the TWFE estimator into a weighted average of all possible $2\times 2$ DiD comparisons between pairs of groups and time periods during which one group enters treatment and the other does not. The terms in his decomposition are not two-period $ATT(g,t)$-type estimators; they aggregate over the relevant pre- and post-periods just like $\beta^{OLS}$ from (5.4) does. They include many comparisons to already-treated units like (5.3). This mechanical decomposition always has positive weights, and they are larger for larger groups and for groups treated closer to the middle of the panel, which have a larger variance of $D_{i,t}$ conditional on the fixed effects. This result shows how TWFE estimators actually function and creates a clear link to the estimation approaches we outlined above for summary parameters. Both are averages of $2\times 2$ DiD terms, but they differ in which comparisons they use and how they aggregate.
Moving from a decomposition of the estimator to a decomposition of the estimand further clarifies why the mechanics of TWFE likely do not identify a desirable parameter, even if parallel trends holds. Unfortunately, in staggered designs, TWFE estimands do not have a clear causal interpretation under PT-type assumptions alone, because they incorporate many already-treated comparisons like the ones in (5.18) (Goodman2021; deChaisemartin2020; Borusyak2023; Imai2021; Strezhnev2018; Sun2021). This will tend to bias $\beta^{twfe}$ away from the sign of $ATT_{\text{avg}}$, and $\beta^{twfe}$ can even have the opposite sign of $ATT_{\text{avg}}$ (baker2022much). It may appear that a more flexible regression specification could solve this problem, but Sun2021 show that a TWFE event study specification suffers from a similar bias when the dynamics of the $ATT(g,t)$’s differ across cohorts. Moreover, the “variance-weighting” feature of OLS means that $\beta^{twfe}$ has non-intuitive weights even when $ATT(g,t)=ATT(g)$.
While TWFE remains common, it has well-understood, potentially serious, and easily remedied problems, and we do not recommend using it. In many cases, especially with many untreated units or minimal treatment effect dynamics, TWFE estimates may be similar to those derived from the theoretically grounded estimators discussed above. The only way to be sure, however, is to estimate both. In that case, it is unclear why a researcher would not report the estimates from a procedure motivated by a desirable target parameter and a credible PT assumption.
@@ 6 Conclusion
    Justify the estimation method. In some DiD designs, estimation is as simple as replacing population expectations with sample means. In others, such as conditional DiD designs, estimation involves choosing econometric techniques (e.g., a regression adjustment, inverse probability weighting, or doubly robust procedure) to map theoretical quantities to estimable sample quantities. Each of these strategies relies on additional modeling restrictions that should be stated clearly and respect the identification assumptions.
@@
    Conduct sensitivity analysis. A clear statement of the identification and estimation assumptions also facilitates a clear statement of what violations of those assumptions might mean. No study is robust to all the ways its assumptions may fail, but a good study should be robust against likely violations of plausible magnitudes. Combine context-specific knowledge about how the assumptions from Step 2 might be violated and by how much with the structure of the estimator from Step 3 to evaluate how much the DiD estimates vary if the key identification assumptions are not exactly true.
@@
    Keep learning. DiD is not the only or even the most plausible research design in all settings; it is just one of many causal inference techniques. If the assumptions required for a DiD analysis appear implausible ex ante or are refuted by evidence or non-robustness in practice, then explore different designs whose assumptions may be more plausible. If existing DiD methods do not provide enough guidance, then use a forward engineering approach to deduce what advances would help.
@@
Although this paper is by no means an exhaustive guide to DiD practice, the eight steps above are a rigorous framework for tackling all the DiD topics that we did not cover. The Appendix briefly discusses DiD methods for (a) treatments that turn on and off over time, (b) continuous and multi-valued treatments, (c) triple differences, (d) distributional parameters, and (e) repeated cross-section or unbalanced panel data. While each of these designs differs from what we covered in the main text, a forward-engineering approach that moves from defining parameters and assumptions, to settling on estimation and inference techniques, to probing robustness, applies equally to all of them. While the specifics of any given DiD analysis may change across research questions, treatment variables, econometric techniques, and data structures, the principles by which one can conduct reliable and transparent causal inference stay the same. [^33] There are other DiD topics of interest that we do not cover, including fuzzy DiD and instrumented DiD designs (DeChaisemartin2018; Miyaji_DiD_IV_2024), nonlinear DiD models (Wooldridge2023; TchetgenTchetgen2024_universalDiD), issues related to few clusters (Roth2023a, Section 5), and situations with multiple treatments (deChaisemartin2023b-intertemporal-treatments; Yanagi2023). We also do not cover some methods that address violations of parallel trends (Freyaldenhoven2019; Arkhangelsky2021_SDiD; Callaway2023; Imbens_Kallus_Mao_2021), or setups that impose treatment timing being as-good-as-random (Athey2022_DiD_design_based; Roth2023c; Arkhangelsky2024_DRTWFE).
@@ A.1 Setups with treatment turning on and off
To tackle this problem from first principles, we need to augment the potential outcomes to reflect the richer notion of treatment sequences. Following Robins1986, let $Y_{i,t}(\textbf{d})$ denote the potential outcome for unit $i$ at time $t$ if this unit received the $T$-dimensional treatment sequence $\textbf{d}\in\{0,1\}^{T}$. For simplicity, let’s say that $T=3$ and that no unit is treated in the first period. In this case, we have four treatment sequences (or histories) that define four potential outcomes for each unit: $Y_{i,t}(0,0,0),Y_{i,t}(0,0,1),Y_{i,t}(0,1,0)$ and $Y_{i,t}(0,1,1)$. We then define treatment groups based on treatment sequences: $G=\textbf{d}_{0}\equiv(0,0,0)$ (never-treated), $G=\textbf{d}_{1}\equiv(0,0,1)$ (treated in the third period), $G=\textbf{d}_{2}\equiv(0,1,1)$ (treated in the second and third period), and $G=\textbf{d}_{3}\equiv(0,1,0)$ (treated only in the second period). In general, we would have as many groups as we have different (realized) treatment sequences. Recall that in a staggered timing design with an absorbing treatment, treatment timing fully characterizes a treatment sequence.
Once potential outcomes and groups are well-defined, one can move to parameters of interest. Similar to the staggered treatment setup in Section 5.2, we consider group-and-time specific ATTs as building blocks, except that groups are now based on more complex treatment sequences. Let 0 denote a T-dimensional vector of zeros. One intuitive building block parameter on which to base a DiD analysis is
@@
Next, one needs to establish identification for the parameters, and propose appropriate estimators and inference procedures. Following similar arguments to those in Section 5.2, a DiD approach to this problem would involve imposing a parallel trends assumption (potentially conditional on covariates) and a no-anticipation assumption to establish that the $ATT(\textbf{d},t)$’s are identified. If each treatment group is sufficiently large, one could proceed in a similar fashion as the staggered setup, comparing average outcome paths for a given sequence to the average outcome path for never-treated (or not-yet-treated) units. One can also aggregate these different $ATT(\textbf{d},t)$ to form different summary parameters.
@@
deChaisemartin2020 imposes a “no-carryover” assumption that implies that past treatments do not affect future outcomes, i.e., that treatment effects in a given period last only during that period. With such an assumption (in addition to parallel trends and a no-anticipation assumption), they propose DiD estimators for an instantaneous average treatment effects parameter by comparing currently treated units with untreated units. Imai2023Matching adopt a similar approach, though they impose a limited-carryover assumption where treatments may last for $\ell$ periods (with $\ell$ specified by the researcher). They then propose estimators for an average treatment effect of switching into treatment in period $t$ among units that experience the policy change in period $t$, and share the same treatment history over the previous $k$ periods; see Liu2022 for a related procedure. Finally, deChaisemartin2023b-intertemporal-treatments avoid making assumptions related to carryover effects and extend the DiD framework in deChaisemartin2020 to allow for treatment effect dynamics. The way they proceed is to first “staggerize” treatment sequences according to first-time of treatment exposure, compute a staggered DiD procedure for this “intention-to-treat” type parameter, and normalize them by a DiD estimate based on the number of treated periods. A potential challenge with deChaisemartin2023b-intertemporal-treatments’s approach is the interpretability of their proposed summary parameter, though we should acknowledge that this is a complex setup.

### v2

@@ 5.2.5 Estimators for staggered designs with covariates
As the discussions in Section 5.2.2 made it clear, we can view the staggered DiD setups as a collection of simpler $2\times 2$ DiD building blocks. A benefit of this interpretation is that, if parallel trends holds only after conditioning on the covariates that determine untreated potential outcome changes, we can easily leverage all the results discussed in Section 4 to identify, estimate and make inference about the $ATT(g,t)$’s, using regression adjustment, inverse probability weighting, or doubly robust methods.
Of course, to proceed in this manner, we would need to adopt conditioned on covariates versions of Assumption PT-GT-Nev, PT-GT-NYT or PT-GT-all, as well as impose an overlap condition. Given that all these are fairly similar to each other, here we state only an extension of Assumption PT-GT-NYT and a strong overlap condition that can be used for all three cases.
@@
For every eventually treated group $g$, not-yet-treated group $g^{\prime}$, time periods $t$ such that $t\geq g$ and $g^{\prime}>t$, and every covariate value $X_{i}$,
@@
For every group $g\in\mathcal{G}$, the conditional (weighted) probability of belonging to a treatment group $g$, given observed covariates $X_{i}$ that are determinants of untreated potential outcome growth, is uniformly bounded away from zero and one. That is, for some $\epsilon>0$ and for every group $g\in\mathcal{G}$, $\epsilon<P_{\omega}[G_{i}=g|X_{i}]<1-\epsilon$.
Using these identifying assumptions and building on the results in Sections 4 and 5.2.2, we can follow the arguments in Callaway2021 and establish that the post-treatment $ATT(g,t)$’s are identified by the regression-adjusted, inverse-probability-weighted, and doubly-robust estimands given by [^31] Wooldridge2021 propose alternative estimators for the $ATT(g,t)$ that incorporate covariates and their interactions with group and time dummies into (5.12). Although Borusyak2023 and deChaisemartin2020 also allow for covariates in their estimation procedure, they affect the outcome levels and not their changes over time. Thus, to some extent, these procedures do not build on conditional parallel trends like Assumption CPT-GT-NYT.
@@
Estimating the $ATT(g,t)$’s follows exactly as in Section 4, and event study aggregations follow from the arguments in Section 5.2.4. Figure 9 reports event study summary estimates incorporating covariates into the Medicaid analysis. Like Figure 8, it suggests that Medicaid expansion had no effect on adult mortality among counties that expanded Medicaid by 2019. Given the point estimates and uniform confidence interval, we can be reasonably confident that the treatment effects are not greater than 6 or less than 11 deaths per 100,000 adults for the 6 year period following Medicaid expansion.
@@
Notes: This figure shows the event-study estimates with staggered treatment timing, using the doubly-robust estimation method from Callaway2021. The outcome variable is the crude mortality rate for adults ages 20-64, and the covariates include the percentage of the county population that is female, the percentage of the county population that is white, the percentage of the county population that is Hispanic, the unemployment rate, the poverty rate, and county-level median income. The point estimate is reported by the circles, and both 95% point-wise (black) and simultaneous (red) confidence intervals are reported with the vertical lines. We also report the simple average of all non-negative event times as a summary of the overall ATT (together with their standard errors and 95% confidence interval). All results use population weights.
@@ 5.3 Limitations of TWFE regressions
Our framework emphasizes building an estimator from $2\times 2$ components, each of which targets a well-defined $ATT$ parameter under a specific parallel trends assumption. The most common estimator for staggered designs, a TWFE regression, comes instead from extending convenient estimation tools that work well in the $2\times 2$ case. A TWFE specification that estimates a summary treatment effect parameter is:
@@
To isolate this issue, consider a setting with two time periods and three groups: a group that enters treatment in the first period ($G_{i}=1$), a group that becomes treated in the second time period ($G_{i}=2$), and a never treated group ($G_{i}=\infty$). This is a staggered design because group 1 and group 2 are treated at different times, but because there are only two time periods, we can re-write the TWFE specification as
$\Delta Y_{i,2}=\Delta\theta_{t}+\beta^{twfe}\Delta D_{i,2}+\Delta e_{i,2}.$
Because $\Delta D_{i,2}$ takes only two values—1 for units whose treatment status increases ($G_{i}=2$), and 0 for units whose treatment status does not change ($G_{i}=1$ and $G_{i}=\infty$)—the TWFE estimand is the following simple comparison of means:
$\beta^{twfe}$ $=$ $\mathop{}\!\mathbb{E}[\Delta Y_{i,2}|\Delta D_{i,2}=1]-\mathop{}\!\mathbb{E}[\Delta Y_{i,2}|\Delta D_{i,2}=0]$ (5.16)
$=$ $\Big{(}\mathop{}\!\mathbb{E}[\Delta Y_{i,2}|G_{i}=2]-\mathop{}\!\mathbb{E}[\Delta Y_{i,2}|G_{i}=\infty]\Big{)}(1-w_{1})+\Big{(}\mathop{}\!\mathbb{E}[\Delta Y_{i,2}|G_{i}=2]-\mathop{}\!\mathbb{E}[\Delta Y_{i,2}|G_{i}=1]\Big{)}w_{1},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak$
where $w_{1}=\frac{p_{1}}{p_{1}+p_{\infty}}$ and $p_{g}=P(G=g)$ is group $g$’s share of units. The TWFE coefficient, in this case, is a weighted average of two DiD terms that use the already-treated units or the never-treated units as comparisons. We have already discussed how under PT between group $G_{i}=2$ and never-treated units, the first term in the $\widehat{\beta}^{twfe}$ decomposition, $\mathop{}\!\mathbb{E}[\Delta Y_{i,2}|G_{i}=2]-\mathop{}\!\mathbb{E}[\Delta Y_{i,2}|G_{i}=\infty]$, equals $ATT(2,2)$. But what about the second term with the already treated comparison group? It turns out that under parallel trends and no-anticipation, if we add and subtract different terms, this type of estimand generally equals a combination of treatment effects for both groups:
$\mathop{}\!\mathbb{E}[\Delta Y_{i,2}|G_{i}=2]-\mathop{}\!\mathbb{E}[\Delta Y_{i,2}|G_{i}=1]$ $=$ $\mathop{}\!\mathbb{E}[Y_{i,2}|G_{i}=2]-\mathop{}\!\mathbb{E}[Y_{i,1}|G_{i}=2]$
$-\Big{(}\mathop{}\!\mathbb{E}[Y_{i,2}|G_{i}=1]-\mathop{}\!\mathbb{E}[Y_{i,1}|G_{i}=1]\Big{)}$
$=$ $\mathop{}\!\mathbb{E}[Y_{i,2}(2)-Y_{i,2}(\infty)|G_{i}=2]+\mathop{}\!\mathbb{E}[Y_{i,2}(\infty)-Y_{i,1}(\infty)|G_{i}=2]$
$-\Big{(}\mathop{}\!\mathbb{E}[Y_{i,2}(1)-Y_{i,2}(\infty)|G_{i}=1]-\mathop{}\!\mathbb{E}[Y_{i,1}(1)-Y_{i,1}(\infty)|G_{i}=1]\Big{)}$
$-\mathop{}\!\mathbb{E}[Y_{i,2}(\infty)-Y_{i,1}(\infty)|G_{i}=1]$
$=$ $ATT(2,2)-\Big{(}ATT(1,2)-ATT(1,1)\Big{)}$
$+\mathop{}\!\mathbb{E}[Y_{i,2}(\infty)-Y_{i,1}(\infty)|G_{i}=2]-\mathop{}\!\mathbb{E}[Y_{i,2}(\infty)-Y_{i,1}(\infty)|G_{i}=1]$
$=$ $ATT(2,2)-\Big{(}ATT(1,2)-ATT(1,1)\Big{)},$
where the first equality follows from the linearity of expectations; in the second, we add and subtract averages of untreated potential outcomes and explore no-anticipation, and the observation rules that $Y_{i,t}(g)$ is observed for units with $G_{i}=g$; in the third equality we rearrange terms; and the last equality follows from a parallel trends for units in group $G=1$ and $G=2$.
@@
$\beta^{twfe}=ATT(2,2)+ATT(1,1)w_{1}-ATT(1,2)w_{1}.$ (5.17)
The $ATT(1,2)$ building block receives negative weight in the overall TWFE estimand unless there are no treatment effect dynamics ($ATT(1,1)=ATT(1,2)$). In this example, the problem is easy to fix: drop the always-treated units and target $ATT(2,2)$. With multiple periods, however, the problem is more complex.
There are two primary results regarding TWFE for general staggered timing designs. One is a decomposition of the TWFE estimator. Goodman2021 expresses the TWFE estimator as a weighted average of all possible $2\times 2$ DiD comparisons between pairs of groups and time periods during which one group enters treatment and the other does not. The terms in his decomposition are not two-period $ATT(g,t)$-type estimators; they aggregate over the relevant pre- and post-periods just as $\beta^{OLS}$ from (5.4) does. They include many comparisons to already-treated units like (5.16). This mechanical decomposition always has strictly positive weights that are larger for larger groups and for groups treated closer to the middle of the panel, which have a larger variance of $D_{i,t}$ conditional on the fixed effects. This result shows how TWFE estimators actually function and creates a clear link to the estimation approaches we outlined above for summary parameters. Both are averages of $2\times 2$ DiD terms, but they differ in which comparisons they use and how they aggregate.
The second result is a decomposition of the TWFE estimand, which clarifies why the TWFE estimator is not guaranteed to identify a desirable parameter in staggered designs, even if parallel trends holds. The reason is their use of already-treated comparisons like the ones in (5.17) (Goodman2021; deChaisemartin2020; Borusyak2023; Imai2021; Strezhnev2018; Sun2021). This will tend to bias $\beta^{twfe}$ away from the sign of $ATT_{\text{avg}}$, and $\beta^{twfe}$ can even have the opposite sign of $ATT_{\text{avg}}$ (baker2022much). It may appear that a more flexible regression specification could solve this problem, but Sun2021 show that a TWFE event study specification suffers from a similar bias when the dynamics of the $ATT(g,t)$’s differ across cohorts. Moreover, the “variance-weighting” feature of OLS means that $\beta^{twfe}$ has non-intuitive weights even when $ATT(g,t)=ATT(g)$.
While TWFE remains common, it has well-understood, potentially serious, and easily remedied problems, and we do not recommend using it. In many cases, especially those with many untreated units or minimal treatment effect dynamics, TWFE estimates may be similar to those derived from the theoretically grounded estimators discussed above. The only way to be sure, however, is to estimate both. In that case, it is unclear why a researcher would not report the estimates from a procedure motivated by a desirable target parameter and a credible PT assumption.
@@ 6 Conclusion
    Determine the appropriate estimation method. In some DiD designs, estimation is as simple as replacing population expectations with sample means. In others, such as conditional DiD designs, estimation involves choosing econometric techniques (e.g., a regression adjustment, inverse probability weighting, or doubly robust procedure) to map theoretical quantities to estimable sample quantities. Each of these strategies relies on additional modeling restrictions that should be stated clearly.
@@
    Conduct sensitivity analysis. A clear statement of the identification and estimation assumptions also facilitates a clear statement of what violations of those assumptions might mean. No study is robust to all the ways its assumptions may fail, but a good study should be robust against likely violations of plausible magnitudes. Combine context-specific knowledge about how the assumptions from Step 2 might be violated, and by how much, with the structure of the estimator from Step 3 to evaluate how much the DiD estimates vary if the key identification assumptions are not exactly true.
@@
    Keep learning. DiD is not the only or the best research design in all settings; it is just one of many causal inference techniques. If the assumptions required for a DiD analysis appear implausible ex-ante or are refuted by evidence or non-robustness in practice, then explore different designs. If existing DiD methods do not provide enough guidance, then use a forward engineering approach to deduce what advances would help.
@@
Although this paper is by no means an exhaustive guide to DiD practice, the eight steps above are a rigorous framework for tackling all the DiD topics that we did not cover. The Appendix briefly discusses DiD methods for (a) treatments that turn on and off over time, (b) continuous and multi-valued treatments, (c) triple differences, (d) distributional parameters, and (e) repeated cross-section or unbalanced panel data. While each of these designs differs from what we covered in the main text, a forward-engineering approach that moves from defining parameters and assumptions, to settling on estimation and inference techniques, to probing robustness, applies equally to all of them. While the specifics of any given DiD analysis may change across research questions, treatment variables, econometric techniques, and data structures, the principles by which one can conduct reliable and transparent causal inference stay the same. [^32] There are other DiD topics of interest that we do not cover, including fuzzy DiD and instrumented DiD designs (DeChaisemartin2018; Miyaji_DiD_IV_2024), nonlinear DiD models (Wooldridge2023; TchetgenTchetgen2024_universalDiD), issues related to few clusters (Roth2023a, Section 5), and situations with multiple treatments (deChaisemartin2023b-intertemporal-treatments; Yanagi2023). We also do not cover some methods that address violations of parallel trends (Freyaldenhoven2019; Arkhangelsky2021_SDiD; Callaway2023; Imbens_Kallus_Mao_2021), nor do we examine setups that impose as-good-as-random treatment timing (Athey2022_DiD_design_based; Roth2023c; Arkhangelsky2024_DRTWFE).
@@ A.1 Setups with treatment turning on and off
To tackle this problem from first principles, we need to augment the potential outcomes to reflect the richer notion of treatment sequences. Following Robins1986, let $Y_{i,t}(\textbf{d})$ denote the potential outcome for unit $i$ at time $t$ if this unit received the $T$-dimensional treatment sequence $\textbf{d}\in\{0,1\}^{T}$. For simplicity, let’s say that $T=3$ and that no unit is treated in the first period. In this case, we have four treatment sequences (or histories), which define four potential outcomes for each unit: $Y_{i,t}(0,0,0),Y_{i,t}(0,0,1),Y_{i,t}(0,1,0)$ and $Y_{i,t}(0,1,1)$. We then define treatment groups by treatment sequences: $G=\textbf{d}_{0}\equiv(0,0,0)$ (never-treated), $G=\textbf{d}_{1}\equiv(0,0,1)$ (treated in the third period), $G=\textbf{d}_{2}\equiv(0,1,1)$ (treated in the second and third period), and $G=\textbf{d}_{3}\equiv(0,1,0)$ (treated only in the second period). In general, we would have as many groups as we have different (realized) treatment sequences. Recall that in a staggered timing design with an absorbing treatment, treatment timing fully characterizes a treatment sequence.
Once potential outcomes and groups are well-defined, one can move to parameters of interest. We proceed similarly to the staggered treatment setup in Section 5.2 and consider group-and-time specific ATTs as building blocks, except that groups are now based on more complex treatment sequences. Let 0 denote a T-dimensional vector of zeros. One intuitive building block parameter on which to base a DiD analysis is
@@
Next, one needs to establish identification for the parameters, and propose appropriate estimators and inference procedures. Following similar arguments to those in Section 5.2, a DiD approach to this problem would involve imposing a parallel trends assumption (potentially conditional on covariates) and a no-anticipation assumption to establish that the $ATT(\textbf{d},t)$’s are identified. If each treatment group is sufficiently large, one could proceed in a similar fashion as the staggered setup, comparing average outcome paths for a given sequence with the average outcome path for never-treated (or not-yet-treated) units. One can also aggregate these different $ATT(\textbf{d},t)$ to form different summary parameters.
@@
deChaisemartin2020 impose a “no-carryover” assumption that implies that past treatments do not affect future outcomes; which is to say that treatment effects in a given period last only during that period. With such an assumption (in addition to parallel trends and a no-anticipation assumption), they propose DiD estimators for an instantaneous average treatment effects parameter by comparing currently treated units with untreated units. Imai2023Matching adopt a similar approach, though they impose a limited-carryover assumption where treatments may last for $\ell$ periods (with $\ell$ specified by the researcher). They then propose estimators for an average treatment effect of switching into treatment in period $t$ among units that experience the policy change in period $t$, and share the same treatment history over the previous $k$ periods; see Liu2022 for a related procedure. Finally, deChaisemartin2023b-intertemporal-treatments avoid making assumptions related to carryover effects and extend the DiD framework in deChaisemartin2020 to allow for treatment effect dynamics. The way they proceed is to first “staggerize” treatment sequences according to first-time of treatment exposure, compute a staggered DiD procedure for this “intention-to-treat” type parameter, and normalize them by a DiD estimate based on the number of treated periods. A potential challenge with deChaisemartin2023b-intertemporal-treatments’s (deChaisemartin2023b-intertemporal-treatments) approach is the interpretability of their proposed summary parameter, though we should acknowledge that this is a complex setup.